# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ V-D2 Network Routing
Throughput |
Tx Delay/jitter |
Error Rate |
SL | UL | RL |
RL Model |
DL Model |
[300] |
Q-Learning |
[301] |
Q-Learning |
[302] |
Distributed Q-Learning |
[303] |
Fuzzy-CNN |
[127] |
CNN |
[304] |
FCNN |
[305] |
ELM |
[306] |
Graph Neural Network (GNN) + FCNN |
[307] |
DDPG |
LSTM |
@@
In wireless networks, depending on the traffic and channel conditions, the network traffics are routed to different paths among different network nodes so that the overall network performance can be improved. In any network with static channel and traffic conditions, this routing problem can be transformed into the well-known shortest path problem and solved by Dijkstra’s algorithm [308]. Here the network nodes can be considered as nodes in the graph and the edges can represent the links between different nodes. The weights of the edges can be defined based on the target network performance metrics like delay, jitter, throughput, packet loss, etc. However, the topology of the real satellite-terrestrial integrated networks (shown in Figure 18) are very complex and dynamic due to hierarchical network architecture and uncertain channel and traffic conditions, respectively. So simple Djikstra’s algorithm cannot be directly applied to meet the performance requirements in these networks.
@@
In real satellite networks, Asynchronous Transfer Mode (ATM) routing is introduced in [309]. The well-known Open Shortest Path First (OSPF) [310] protocol-based Internet Protocol (IP) routing is also adapted to the dynamic satellite environment in [311]. However, the dynamics of satellite-terrestrial networks are very different from traditional terrestrial networks due to their highly dynamic network topology, link status, and traffic conditions. Depending on the instantaneous network topology, a static-dynamic combined routing scheme is considered in [312]. An ant-colony-based optimization (ACO) framework is considered in [313]. In [314], a Kalman filter-based Wolf Colony Optimization algorithm is used to solve the local optimal solution issue in [313]. An improved ACO framework is considered to find out the optimal set of links with multiple network constraints in [315]. A Coordinate Graph (CG) model-based network routing approach for three-dimensional TNTN is considered in [316]. In [317], minimum flow maximum residual path-based network flow algorithm is used to find out the optimal network routing path for satellites. In [318], a 3-dimensional network mapping using hyperbolic geometry is considered for integrated satellite-terrestrial networks.
@@
To cope with the dynamic environment in integrated satellite-terrestrial networks, different DL architectures are proposed in the literature. In [303], fuzzy logic is used to evaluate task requirements to improve the CNN output for optimal path allocation. The network routing optimization problem can be put into an RL framework. In [300], a speed-up Q-learning algorithm is used to find out the optimal routing strategy for TNTN. A similar Q-learning-based RL framework is also considered to solve the routing problem for LEO satellites in [301]. To tackle the complexity issue, a DRL framework is used to generate optimal routing strategies in TNTN [307] and LEO satellite networks [301]. FCNN [304], CNN [127], etc. architectures are used to solve the routing problem in a supervised manner. Other ML frameworks like GNN [306] and ELM [305] are also considered to solve the routing problem in NTNs. In table X, we summarize the AI approaches for network routing in TNTN. Recursive NN architectures need to be also explored for capturing the temporal behavior in network routing decisions. Furthermore, the channels are extremely dynamic and time-varying in the case of TNTNs; the channel conditions can be also considered in the learning criteria of RL frameworks.
@@ V-D3 Traffic Prediction
Traffic prediction is very critical in modern communication systems to ensure high-speed low latency communications. Particularly, in NTNs, accurate traffic prediction is extremely crucial due to the highly dynamic network topology as well as diverse user requirements. At its core, traffic prediction involves forecasting future network traffic based on past usage patterns. Conventional approaches such as Auto Regressive Moving Average (ARMA), Auto Regressive Integrated Moving Average (ARIMA) [319, 320], etc. are typically used for these predictions. However, the DL approaches have emerged as more effective alternatives by providing improved performances in recent times due to their inherent capability to capture spatial and temporal correlations.
In [321], Radial Basis Functions (RBF) neural network-based short-term traffic flow forecasting is proposed. In [322], an LSTM-based architecture is utilized for traffic prediction due to its temporal characteristics handling capability where the attention mechanism is used to balance the effect of inputs on outputs properly. The RNN architectures suffer from gradient explosion issues. To overcome this issue, GRU architectures are explored for traffic prediction in [323, 324, 325]. In [323], the transfer learning approach and particle filter online training algorithm are combined to address the lack of online training data and reduce the training time complexity. In [324], GNNs are used to extract the spatial features of the satellite network traffic from the input network topology, which is later used as an input to a GRU network for traffic prediction. In [325], on top of the attention mechanism and GRU models, PSO is used to obtain the best set of hyperparameters for the network.
@@ VI AI-NTN Integration: Current Status
In the previous section, we discuss how AI can be beneficial for us in resolving potential NTN issues for the next-generation 6G networks. In this section, how AI can be applied to real systems to resolve the challenges associated with NTNs. We begin our discussion by discussing the current ML testbeds for satellites. Then we discuss how AI can be potentially applied to future 6G networks by utilizing the RAN Intelligent Controller (RIC) embedded in the Open Radio Access Network (O-RAN) framework [326] to overcome the inflexibility of the monolithic cellular networks. Finally, we provide a discussion on current research efforts toward realizing the Software Defined Radio (SDR) based 5G-NTN platform development in the O-RAN framework.
@@ VI-A ML Testbeds for Satellite Networks
MultI-layer awaRe SDN-based testbed for SAtellite-Terrestrial networks (MIRSAT) testbed [327] provides a Software Defined Network (SDN) based experimentation platform for testing network slicing algorithms on NGSO constellations. The European Space Agency (ESA) has numerous completed projects focusing on the applicability of AI techniques in satellite networks such as MLSAT [328] and SATAI [329]. There are also several other ongoing projects of ESA focusing on AI-satellite issues like AI integrated 5G-Satellite testbed [330], AI-based interference detection [331], AI-based signal processing [332], etc. All these testbeds show promises for AI to be an integral part of future satellite-terrestrial integrated networks.
@@ VI-B AI-NTN Integration through O-RAN-based RIC
Traditional 5G networks with little or no reconfiguration capabilities suffer from a wide variety of challenges to satisfy the heterogeneity and variability of the networks and meet the strict application requirements [333]. Even though there has been a significant amount of research on addressing different issues in 5G cellular networks, an open interface for the deployment of AI algorithms is required. O-RAN offers a general framework for the deployment of AI algorithms in 5G-Advanced networks [334]. It achieves this by facilitating an open interface that enables the exchange of network KPIs and control information between the RAN and the AI controller. This integration allows for the implementation of a closed-loop control framework for the RAN using different AI approaches [335, 336]. As a potential integral part of future 6G networks, NTNs are expected to be deployed in the O-RAN framework to leverage AI capabilities effectively.
@@
The non-real-time which does not need to consider latency requirements, is expected to be deployed on the ground considering power, onboard capability, and mobility constraints. However, the near real-time RIC needs to be close to DUs to provide near real-time control functionalities which provide two different options for its deployments with corresponding pros and cons. The near real-time RIC should be on the ground when only RU is in the air, whereas it should be also in the air in the other two cases. There is a clear trade-off between the latency and power, mobility, and onboard capability constraints. If the near-real-time RIC is in the air, the latency for control commands will be low, whereas the cost will be high for hosting it in the air. In Figure 20, the potential framework for AI-Enabled NTN deployment in O-RAN framework as specified in [326] is illustrated. Depending on the deployment scenarios of the near-real-time RIC, the xApps can be also deployed also in the air or on the ground, and so do the AI control algorithms.
@@ VI-C Current Research Efforts for AI-NTN Integration through O-RAN-based RIC
To facilitate the integration of NTN into 6G networks, there have been already some technical advancements and experimental research works towards developing real prototypes for testing and evaluation of proof-of-concept methods. OpenAirInterface (OAI) is an open-source 3GPP compliant SDR-based protocol stacks that are widely used across the research community for experimentation with 5G networks [337]. As specified in the O-RAN framework, OAI protocol stacks adopt the notion of RIC by enabling service-oriented controllers using an efficient Software Development Kit, called Flexible RIC (FlexRIC) [338]. This RIC provides an interface for applying AI algorithms in order to optimize the network performances through xApps as discussed before. This enables us to perform experiments for testing diverse AI approaches for optimizing the performances of real 5G networks.
OAI has been adopted for developing experimental prototypes with 5G-NTN adaptations due to its efficient and flexible design and structure [339]. Currently, there are several research projects on 5G-NTN that are leveraging OAI protocol stacks to perform experiments with NTN adaptations for both in-lab validation and over-the-satellite testing. 5G AgiLe and flexible integration of SaTellite And cellulaR (5G-ALLSTAR) [340] and 5G New Radio EMUlation over SATellite (5G-EmuSat) [341] project developed a 5G-NTN platform with necessary PHY and MAC layer 5G-NR adaptations on top of OAI 5G protocol stacks and a satellite-channel emulator for in-lab validation. 5G-EmuSat even has also demonstrated its over-the-satellite capability by having direct access to a UE using a satellite channel. 5G Space Communications lab also has performed in-lab validation experiments extending OAI-4G protocol stacks for NTN along with ISL implementation using SDR [342]. Two current ongoing projects focusing on GEO and LEO satellites, named 5G-GOA [343] and 5G-LEO [344] respectively, are currently working on implementing necessary 3GPP NTN adaptations extending from 5G protocol stacks of OAI. Even though current implementations are mostly for demonstration purposes, integration of NTN into OAI 5G protocol stacks paves the way for deploying AI algorithms through xApps in the future.
@@ VII-A Limited Onboard Capability
Advanced AI applications necessitate specialized AI-capable embedded chipsets developed by leading technology companies, including NVIDIA, AMD, Intel, and Qualcomm [345]. Typically, these chipsets feature on-chip accelerators like CUDA and Tensor, enabling highly efficient parallel processing of intricate data tasks, especially those involving extensive matrix operations as utilized in DL methodologies. The performance of these AI algorithms highly depends on the availability of the computational resources and data processing capabilities of the AI hardware blocks. More computational capability typically means more power consumption as well as more physical space which increases the overall maintenance cost for the non-terrestrial platforms. Most non-terrestrial platforms such as satellites have a very limited amount of computational resources due to cost optimization. The efficiency of computing devices, particularly those used for onboard purposes, is often quantified by evaluating the ratio of computational power to the product of power consumption, total mass, and associated cost. This metric, highlighted in [346], necessitates a notably high value for onboard computing devices in satellites. All these AI-capable chipsets must undergo meticulous design and rigorous testing procedures to ensure they satisfy the minimum computing efficiency standards mandated for space platforms. A lot more advancements in miniaturization and power efficiency are necessary to ensure the adaption of AI models and algorithms into non-terrestrial platforms with adequate onboard capabilities as well.
@@ VII-D Security Aspects
Applying AI in NTNs introduces more vulnerability to various security attacks by introducing new attack surfaces and less transparency. Adversarial attacks, data poisoning, and model evasion involving the manipulation of input data to AI models can cause degradation in network performance and reliability [347]. Since this input data is gathered from the NTN environment, featuring various attack interfaces for potential attackers, there is a risk that the training data used for AI models could be compromised by these attackers. Denial-of-Service (DoS) attacks can cause interruptions in crucial network operations by overwhelming the network with too much resource consumption [348]. As an illustrative example within the context of NTNs, an attacker could conceivably gain access to one of the network slices. They could then exploit this access to excessively consume network resources by setting up extreme requirements, potentially leading to network congestion and subsequently causing a decline in overall network performance. As discussed before, the AI controller and the network environment, especially in the online setup, needs to exchange control information during training of the models. While carrying the control information between the AI controller and the network, an attacker can intercept and possibly modify this information, which is known as a Man-in-the-Middle (MiM) attack [349]. This can often result in a degradation in network performances due to the compromised control information. Therefore, A security-constrained framework for deploying AI models needs to be designed carefully in such a way that they can detect and mitigate these attacks while maintaining the overall network performance.
@@ VII-E Environmental Conditions
NTN platforms, especially satellites are generally deployed in pretty hostile environments with extreme radiation, extreme temperatures, and other extreme environmental conditions. The computational hardware for AI models is very susceptible to radiation as they are built on customized circuitry [346]. The satellites need to consider both single-event effects caused by ionizing particles [350] as well as the effects due to long-term radiation [351]. These effects may result in bit flips both in registers and memories introducing errors in the control logic of their hosting hardware platforms. Also, in space, these circuitry elements need to withstand extreme temperatures for a long period of time [352]. To ensure the proper performance of these models, the hosting hardware components are required to be more advanced and have rigorous testing to ensure extreme environmental tolerance, which increases the cost of the satellite operation.

### v4

@@ V-D2 Network Routing
| Throughput | Tx Delay/jitter | Error Rate | SL | UL | RL | RL Model | DL Model | ||
| [ntn-rt-rl1] | Q-Learning | ||||||||
| [leo-rt-rl1] | Q-Learning | ||||||||
| [leo-distributed-rt-rl2] | Distributed Q-Learning | ||||||||
| [ntn-rt-fuzzy-cnn] | Fuzzy-CNN | ||||||||
| [IWC_AI_NTN_2] | CNN | ||||||||
| [ntn-rt-dl2] | FCNN | ||||||||
| [ntn-rt-dist-ml1] | ELM | ||||||||
| [leo-rt-gnn] | Graph Neural Network (GNN) + FCNN | ||||||||
| [ntn-rt-drl1] | DDPG | LSTM | |||||||
@@
In wireless networks, depending on the traffic and channel conditions, the network traffics are routed to different paths among different network nodes so that the overall network performance can be improved. In any network with static channel and traffic conditions, this routing problem can be transformed into the well-known shortest path problem and solved by Dijkstra’s algorithm [rt-dj]. Here the network nodes can be considered as nodes in the graph and the edges can represent the links between different nodes. The weights of the edges can be defined based on the target network performance metrics like delay, jitter, throughput, packet loss, etc. However, the topology of the real satellite-terrestrial integrated networks (shown in Figure 18) are very complex and dynamic due to hierarchical network architecture and uncertain channel and traffic conditions, respectively. So simple Djikstra’s algorithm cannot be directly applied to meet the performance requirements in these networks.
@@
In real satellite networks, Asynchronous Transfer Mode (ATM) routing is introduced in [ntn-rt-atm]. The well-known Open Shortest Path First (OSPF) [ospf] protocol-based Internet Protocol (IP) routing is also adapted to the dynamic satellite environment in [ntn-rt-ip]. However, the dynamics of satellite-terrestrial networks are very different from traditional terrestrial networks due to their highly dynamic network topology, link status, and traffic conditions. Depending on the instantaneous network topology, a static-dynamic combined routing scheme is considered in [ntn-rt-adapt]. An ant-colony-based optimization (ACO) framework is considered in [ntn-rt-aco]. In [ntn-rt-kfwca], a Kalman filter-based Wolf Colony Optimization algorithm is used to solve the local optimal solution issue in [ntn-rt-aco]. An improved ACO framework is considered to find out the optimal set of links with multiple network constraints in [ntn-rt-opt]. A Coordinate Graph (CG) model-based network routing approach for three-dimensional TNTN is considered in [ntn-rt-tvgraph]. In [ntn-rt-set], minimum flow maximum residual path-based network flow algorithm is used to find out the optimal network routing path for satellites. In [ntn-rt-hygeo], a 3-dimensional network mapping using hyperbolic geometry is considered for integrated satellite-terrestrial networks.
@@
To cope with the dynamic environment in integrated satellite-terrestrial networks, different DL architectures are proposed in the literature. In [ntn-rt-fuzzy-cnn], fuzzy logic is used to evaluate task requirements to improve the CNN output for optimal path allocation. The network routing optimization problem can be put into an RL framework. In [ntn-rt-rl1], a speed-up Q-learning algorithm is used to find out the optimal routing strategy for TNTN. A similar Q-learning-based RL framework is also considered to solve the routing problem for LEO satellites in [leo-rt-rl1]. To tackle the complexity issue, a DRL framework is used to generate optimal routing strategies in TNTN [ntn-rt-drl1] and LEO satellite networks [leo-rt-rl1]. FCNN [ntn-rt-dl2], CNN [IWC_AI_NTN_2], etc. architectures are used to solve the routing problem in a supervised manner. Other ML frameworks like GNN [leo-rt-gnn] and ELM [ntn-rt-dist-ml1] are also considered to solve the routing problem in NTNs. In table X, we summarize the AI approaches for network routing in TNTN. Recursive NN architectures need to be also explored for capturing the temporal behavior in network routing decisions. Furthermore, the channels are extremely dynamic and time-varying in the case of TNTNs; the channel conditions can be also considered in the learning criteria of RL frameworks.
@@ V-D3 Traffic Prediction
Traffic prediction is very critical in modern communication systems to ensure high-speed low latency communications. Particularly, in NTNs, accurate traffic prediction is extremely crucial due to the highly dynamic network topology as well as diverse user requirements. At its core, traffic prediction involves forecasting future network traffic based on past usage patterns. Conventional approaches such as Auto Regressive Moving Average (ARMA), Auto Regressive Integrated Moving Average (ARIMA) [tf-arima, tf-arima-2], etc. are typically used for these predictions. However, the DL approaches have emerged as more effective alternatives by providing improved performances in recent times due to their inherent capability to capture spatial and temporal correlations.
In [ntn-tf1], Radial Basis Functions (RBF) neural network-based short-term traffic flow forecasting is proposed. In [ntn-tp-lstm], an LSTM-based architecture is utilized for traffic prediction due to its temporal characteristics handling capability where the attention mechanism is used to balance the effect of inputs on outputs properly. The RNN architectures suffer from gradient explosion issues. To overcome this issue, GRU architectures are explored for traffic prediction in [ntn-tp-gru, ntn-tp-gcn_gru, ntn-tp-mod-gru]. In [ntn-tp-gru], the transfer learning approach and particle filter online training algorithm are combined to address the lack of online training data and reduce the training time complexity. In [ntn-tp-gcn_gru], GNNs are used to extract the spatial features of the satellite network traffic from the input network topology, which is later used as an input to a GRU network for traffic prediction. In [ntn-tp-mod-gru], on top of the attention mechanism and GRU models, PSO is used to obtain the best set of hyperparameters for the network.
@@ VI AI-NTN Integration: Current Status
In the previous section, we discuss how AI can be beneficial for us in resolving potential NTN issues for the next-generation 6G networks. In this section, how AI can be applied to real systems to resolve the challenges associated with NTNs. We begin our discussion by discussing the current ML testbeds for satellites. Then we discuss how AI can be potentially applied to future 6G networks by utilizing the RAN Intelligent Controller (RIC) embedded in the Open Radio Access Network (O-RAN) framework [oran] to overcome the inflexibility of the monolithic cellular networks. Finally, we provide a discussion on current research efforts toward realizing the Software Defined Radio (SDR) based 5G-NTN platform development in the O-RAN framework.
@@ VI-A ML Testbeds for Satellite Networks
MultI-layer awaRe SDN-based testbed for SAtellite-Terrestrial networks (MIRSAT) testbed [mirsat] provides a Software Defined Network (SDN) based experimentation platform for testing network slicing algorithms on NGSO constellations. The European Space Agency (ESA) has numerous completed projects focusing on the applicability of AI techniques in satellite networks such as MLSAT [mlsat] and SATAI [satai]. There are also several other ongoing projects of ESA focusing on AI-satellite issues like AI integrated 5G-Satellite testbed [anchor], AI-based interference detection [skymon], AI-based signal processing [spaice], etc. All these testbeds show promises for AI to be an integral part of future satellite-terrestrial integrated networks.
@@ VI-B AI-NTN Integration through O-RAN-based RIC
Traditional 5G networks with little or no reconfiguration capabilities suffer from a wide variety of challenges to satisfy the heterogeneity and variability of the networks and meet the strict application requirements [oran-2]. Even though there has been a significant amount of research on addressing different issues in 5G cellular networks, an open interface for the deployment of AI algorithms is required. O-RAN offers a general framework for the deployment of AI algorithms in 5G-Advanced networks [oran-3]. It achieves this by facilitating an open interface that enables the exchange of network KPIs and control information between the RAN and the AI controller. This integration allows for the implementation of a closed-loop control framework for the RAN using different AI approaches [oai-ml-1, oai-ml-2]. As a potential integral part of future 6G networks, NTNs are expected to be deployed in the O-RAN framework to leverage AI capabilities effectively.
@@
The non-real-time which does not need to consider latency requirements, is expected to be deployed on the ground considering power, onboard capability, and mobility constraints. However, the near real-time RIC needs to be close to DUs to provide near real-time control functionalities which provide two different options for its deployments with corresponding pros and cons. The near real-time RIC should be on the ground when only RU is in the air, whereas it should be also in the air in the other two cases. There is a clear trade-off between the latency and power, mobility, and onboard capability constraints. If the near-real-time RIC is in the air, the latency for control commands will be low, whereas the cost will be high for hosting it in the air. In Figure 20, the potential framework for AI-Enabled NTN deployment in O-RAN framework as specified in [oran] is illustrated. Depending on the deployment scenarios of the near-real-time RIC, the xApps can be also deployed also in the air or on the ground, and so do the AI control algorithms.
@@ VI-C Current Research Efforts for AI-NTN Integration through O-RAN-based RIC
To facilitate the integration of NTN into 6G networks, there have been already some technical advancements and experimental research works towards developing real prototypes for testing and evaluation of proof-of-concept methods. OpenAirInterface (OAI) is an open-source 3GPP compliant SDR-based protocol stacks that are widely used across the research community for experimentation with 5G networks [oai-5g]. As specified in the O-RAN framework, OAI protocol stacks adopt the notion of RIC by enabling service-oriented controllers using an efficient Software Development Kit, called Flexible RIC (FlexRIC) [flexric]. This RIC provides an interface for applying AI algorithms in order to optimize the network performances through xApps as discussed before. This enables us to perform experiments for testing diverse AI approaches for optimizing the performances of real 5G networks.
OAI has been adopted for developing experimental prototypes with 5G-NTN adaptations due to its efficient and flexible design and structure [oai-5g-ntn]. Currently, there are several research projects on 5G-NTN that are leveraging OAI protocol stacks to perform experiments with NTN adaptations for both in-lab validation and over-the-satellite testing. 5G AgiLe and flexible integration of SaTellite And cellulaR (5G-ALLSTAR) [5g-allstar] and 5G New Radio EMUlation over SATellite (5G-EmuSat) [5g-emusat] project developed a 5G-NTN platform with necessary PHY and MAC layer 5G-NR adaptations on top of OAI 5G protocol stacks and a satellite-channel emulator for in-lab validation. 5G-EmuSat even has also demonstrated its over-the-satellite capability by having direct access to a UE using a satellite channel. 5G Space Communications lab also has performed in-lab validation experiments extending OAI-4G protocol stacks for NTN along with ISL implementation using SDR [5g-space-com]. Two current ongoing projects focusing on GEO and LEO satellites, named 5G-GOA [5g-goa] and 5G-LEO [5g-leo] respectively, are currently working on implementing necessary 3GPP NTN adaptations extending from 5G protocol stacks of OAI. Even though current implementations are mostly for demonstration purposes, integration of NTN into OAI 5G protocol stacks paves the way for deploying AI algorithms through xApps in the future.
@@ VII-A Limited Onboard Capability
Advanced AI applications necessitate specialized AI-capable embedded chipsets developed by leading technology companies, including NVIDIA, AMD, Intel, and Qualcomm [ai-chip]. Typically, these chipsets feature on-chip accelerators like CUDA and Tensor, enabling highly efficient parallel processing of intricate data tasks, especially those involving extensive matrix operations as utilized in DL methodologies. The performance of these AI algorithms highly depends on the availability of the computational resources and data processing capabilities of the AI hardware blocks. More computational capability typically means more power consumption as well as more physical space which increases the overall maintenance cost for the non-terrestrial platforms. Most non-terrestrial platforms such as satellites have a very limited amount of computational resources due to cost optimization. The efficiency of computing devices, particularly those used for onboard purposes, is often quantified by evaluating the ratio of computational power to the product of power consumption, total mass, and associated cost. This metric, highlighted in [ai-chal], necessitates a notably high value for onboard computing devices in satellites. All these AI-capable chipsets must undergo meticulous design and rigorous testing procedures to ensure they satisfy the minimum computing efficiency standards mandated for space platforms. A lot more advancements in miniaturization and power efficiency are necessary to ensure the adaption of AI models and algorithms into non-terrestrial platforms with adequate onboard capabilities as well.
@@ VII-D Security Aspects
Applying AI in NTNs introduces more vulnerability to various security attacks by introducing new attack surfaces and less transparency. Adversarial attacks, data poisoning, and model evasion involving the manipulation of input data to AI models can cause degradation in network performance and reliability [ai-chal-adv]. Since this input data is gathered from the NTN environment, featuring various attack interfaces for potential attackers, there is a risk that the training data used for AI models could be compromised by these attackers. Denial-of-Service (DoS) attacks can cause interruptions in crucial network operations by overwhelming the network with too much resource consumption [ai-chal-dos]. As an illustrative example within the context of NTNs, an attacker could conceivably gain access to one of the network slices. They could then exploit this access to excessively consume network resources by setting up extreme requirements, potentially leading to network congestion and subsequently causing a decline in overall network performance. As discussed before, the AI controller and the network environment, especially in the online setup, needs to exchange control information during training of the models. While carrying the control information between the AI controller and the network, an attacker can intercept and possibly modify this information, which is known as a Man-in-the-Middle (MiM) attack [ai-chal-mim]. This can often result in a degradation in network performances due to the compromised control information. Therefore, A security-constrained framework for deploying AI models needs to be designed carefully in such a way that they can detect and mitigate these attacks while maintaining the overall network performance.
@@ VII-E Environmental Conditions
NTN platforms, especially satellites are generally deployed in pretty hostile environments with extreme radiation, extreme temperatures, and other extreme environmental conditions. The computational hardware for AI models is very susceptible to radiation as they are built on customized circuitry [ai-chal]. The satellites need to consider both single-event effects caused by ionizing particles [ai-chal-ion] as well as the effects due to long-term radiation [ai-chal-rad]. These effects may result in bit flips both in registers and memories introducing errors in the control logic of their hosting hardware platforms. Also, in space, these circuitry elements need to withstand extreme temperatures for a long period of time [ai-chal-temp]. To ensure the proper performance of these models, the hosting hardware components are required to be more advanced and have rigorous testing to ensure extreme environmental tolerance, which increases the cost of the satellite operation.